If 17.3 g of Li is completely reacted with zinc chloride, how many moles of zinc will be produced.
1
Expert's answer
2020-04-27T01:46:00-0400

Convert mass of Li to moles

n(Li) = m/M = 17.3g/7g/mol = 2.47 mol

2Li + ZnCl2 = 2LiCl + Zn

According to the equation 2mol of Li produces 1 mol of Zn, so

n(Zn) = 2.47/2 = 1.235 mol